OpenStudy (anonymous):

The average of 2 numbers is 17. One of those numbers were 22, what was the other number?

OpenStudy (anonymous):

To find average of two numbers, call numbers a and b\[\frac{ a+b }{ 2 }=average\]

OpenStudy (anonymous):

yes i know that. That does not answer my question though

OpenStudy (anonymous):

\[\frac{ x+22 }{ 2 } = 17\] Multiply both sides by 2 x+22 = 34 -22 -22 x=12

OpenStudy (anonymous):

Thanks!